OpenStudy (anonymous):

how do you convert 185in.=____ ft=_____ in.

OpenStudy (anonymous):

is it 15ft 180 inches

OpenStudy (secret66):

There are 12 inches in a foot...

OpenStudy (anonymous):

15.4167 ft

OpenStudy (anonymous):

nooooooooo ;D since 1 in = 0.0833333 ft

OpenStudy (anonymous):

just multiple 0.0833333 ft to 185 inches

OpenStudy (anonymous):

153.55 annas

OpenStudy (secret66):

Ehmm....I am confused now >.<

OpenStudy (anonymous):

thx redbullz

OpenStudy (secret66):

\[\Large \color{white}{G} \Large \color{White}{o} \Large \color{White} {t} \Large \color{White} {\space} \Large \color{White} {I} \Large \color{White} {\space} \Large \color{white} {t} \Large \color{White} {!!} \]

OpenStudy (unklerhaukus):

\[\quad1[\text{ft}]=12[\text{in}]\\ \frac{1[\text{ft}]}{12[\text{in}]}=\frac{12[\text{in}]}{12[\text{in}]}\\ \frac{1[\text{ft}]}{12[\text{in}]}=1\] \[\quad185[\text{in}]=185[\text{in}]\times1\\ \qquad\qquad=185[\text{in}]\times\frac{1[\text{ft}]}{12[\text{in}]}\\ \qquad\qquad=\frac{185\times1}{12}\frac{[\cancel{\text{in}]}[\text{ft}]}{[\cancel{\text{in}]}}\\\qquad\qquad=\frac{185\times1}{12}[\text{ft}]\]

OpenStudy (unklerhaukus):

185[in] = 180[in] +5[in] 185/12[ft] = 180/12[ft] +5[in]